Question 25 An intangible asset with an estimated useful life of 30 years was acquired on January 1, 2007, for $540,000. On January 1, 2017, a review was made of intangible assets and their expected service lives, and it was determined that this asset had an estimated useful life of 30 more years from the date of the review. What is the amount of amortization for this intangible in 2017? Amount of amortization
Answer #1

Amortization per year = 540,000/30 years = 18,000 per year

Book value after 10 years (Jan 1,2017)

= 540,000 - (18,000*10)

= 360,000

Amortization for 2017 = 360,000/30 years

= 12,000